If two dice are rolled, find the probability of getting a sum of 9.

Enter your answer as a fraction: example 1/5.

Reduce fraction to simplest form.

____________________.
Answer #1

total number of ways=6*6=36

number of ways of  getting a sum of 9= 4 ways

{(3,6),(4,5),(5,4),(6,3)}

the probability of getting a sum of 9=number of ways of  getting a sum of 9/total number of ways

=4/36

=1/9

the above is answer..